[figure: 9.7 cm]

NON-FERMI LIQUID BEHAVIOR IN CeRhSn

A. Ślebarski^{a} P. -C. Ho^{b}, V. S. Zapf^{b}, N. A. Frederick^{b}, M. B. Maple^{b},

A. Jezierski^{c} aInstitute of Physics, University of Silesia, 40-007 Katowice bDepartment of Physics and Institute for Pure and Applied Physical Sciences, University of California, San Diego, La Jolla, CA 92093-0360 cInstitute of Molecular Physics, Polish Academy of Sciences, 60-179 Poznań

Electrical resistivity ρ, magnetic susceptibility χ, specific heat C, X-ray photoelectron spectroscopy (XPS), and the calculations of the electronic structure were made on the compound CeRhSn. The data demonstrate the scaling behavior of the non-Fermi liquid; C(T)/T ∝ χ ∝ T^{−1+λ}, and Δρ ∝ T. We argue that the non-Fermi liquid behavior observed in CeRhSn can be due to the existance of a Griffith’s phase close to a quantum critical point.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : A. Ślebarski

Address for correspondence : Institute of Physics, University of Silesia, 40-007 Katowice

Email address : slebar@physics.ucsd.edu